10, A cylindrical copper wire heater (resistivity of 1.7x10 Ωm) has a radius of 0.5 mm and a length of 10 m. What is the current on this wire when it is connected to a voltage source of 120 V? (a) (b) (c) (d) 2.8 A 65 A 52 A 124 A-, = 0,2302 天心ー。ソー 520A. 1.120 11. What is the equivalent resistance of the circuit between the points A & B? 2 6 2Ω (d) 9Ω 3Ω 12. What is the current through the 8 Ω resistor in the circuit in problem 11 if a 12V battery is connected between points A & B? (a) 1.5A (b) 2A 0.75A 1.0A Te 오!
Answer #1

10)

We know that,

R = rho L/A

R = 1.7 x 10^-8 x 10/(3.14 x 0.0005^2) = 0.22 Ohm

We know that, V = IR=> I = V/R

I = 120/0.23 = 520 A

I = 521 A

(None of the answer given is correct)

-----------------

11)(b) 6 Ohm

The last three are in series,

R1 = 1 + 2 + 3 = 6

R2 = 6 x 3/(6 + 3) = 2

R3 = 2 + 6 = 8

R4 = 8 x 8/(8 + 8) = 4

Req = 4 + 2 = 6 ohm

------------------------------

12)(d)1 A

Current in the circuit is:

I = V/R = 12/6 = 2 A

drop across 2 ohm is:

I2 = 2 x 2 = 4 A

drop across parallel = 12-4 = 8 Ohm

I8 = 8/8 = 1 A